The height reached by a particle in time thrown upward with a speed u is given by h=ut - (1/2)g(t*t) , where g is a constant . Find the time taken in reaching the maximum height.

The  height reached by a particle in time thrown upward with a speed u is given by  h=ut - (1/2)g(t*t) , where g is a constant .   


Find the time taken in reaching the maximum height.

Grade:11

1 Answers

Prabodh Yadav
13 years ago

differentiate this eqn(h) wrt t,i.e.,find dh/dt to get the general expression for velocity(varying with t).Now at max height dh/dt=0 therefore put dh/dt=0 in dh/dt.solve and get t!!

dh/dt=u-gt

now at max height

dh/dt=0

thus,u-gt=0

=>t=u/g